OpenStudy (anonymous):

Let f(x) = 3x. Find f( 2) Please show me all the steps, i'm stumped

OpenStudy (aravindg):

6

OpenStudy (aravindg):

f(x)=3*2=6

OpenStudy (amistre64):

f(x) is notation for the function has a value when x is .... in this case 2

OpenStudy (amistre64):

f(x) = 3x f(2) = 3(2)

OpenStudy (amistre64):

f(x) = 3*2 is bad form ....

OpenStudy (amistre64):

we can use this notation to indicate an ordered pair, like a point on a graph: (x,(f(x))) ; when x = 2 we have: (2,f(2)) = (2,6)

OpenStudy (amistre64):

i got a little to () happy on that one lol

OpenStudy (akshay_budhkar):

LOL amistre!!! i like the way you tell things n explain them hehe

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome :)